OpenStudy (anonymous):

Help no direct answers plz explain Skylar plans to use $3400 to open a savings account with an annual interest rate of 1.15%. How much more interest will he earn over 13 years if he chooses a compound interest account that compounds interest quarterly instead annually? Round your answer to the nearest cent. interest compounded annually: A = P (1 + r)t interest compounded quarterly: A = P (1+r/4)4t

OpenStudy (anonymous):

@Zeronknight @LoveIt @Hayhayz @OswinOswald @heacain

OpenStudy (anonymous):

@jhonyy9

Zeronknight (zeronknight):

I don't know sorry. But l'll tag some people. @rebeccaxhawaii @mathmale

OpenStudy (misssmartiez):

Plug it in. P = 3400 R = 1.15% Time = 13

OpenStudy (misssmartiez):

interest compounded annually: A = 3,400 (1 + 1.15% (put this to decimal if using calculator) )^13 interest compounded quarterly: A = 3,400 (1+1.15%/4)4^13 (is it even exponent? Check this, sometimes then you copy it, it doesn't put it as exponent form. If not, just block out the '^'. USE PEMDAS),.

OpenStudy (anonymous):

I still don't get it

OpenStudy (aajc10):

39.1

OpenStudy (aajc10):

are you still their

OpenStudy (misssmartiez):

PEMDAS: 1: P or ( - ) 1.15 / 100 = 0.0115. 0.0115 + 1 = 1.0115 2: exponent. (is it ^t or regular t?)

OpenStudy (misssmartiez):

Is it exponent T or just plain T. You need to tell me. I don't want to give you the incorrect formula.

rebeccaxhawaii (rebeccaxhawaii):

yes plug those things in like Ms. Smart is telling you then subtract the difference the information is given plug it into the fomulas

OpenStudy (misssmartiez):

^ Thank you Rebecca.

rebeccaxhawaii (rebeccaxhawaii):

no.. thank you. great work ! c:

OpenStudy (aajc10):

wats up

OpenStudy (aajc10):

it is c

OpenStudy (anonymous):

do you still need help